### v2

# Examining Racial Stereotypes in YouTube Autocomplete Suggestions

Abstract
Autocomplete is a popular search feature that predicts queries based on user input and guides users to a set of potentially relevant suggestions. In this study, we examine how YouTube autocompletes serve as an information source for users exploring information about race. We perform an algorithm output audit of autocomplete suggestions for input queries about four racial groups and examine the stereotypes they embody. Using critical discourse analysis, we identify five major sociocultural contexts in which racial biases manifest—Appearance, Ability, Culture, Social Equity, and Manner. Our results show evidence of aggregated discrimination and interracial tensions in the autocompletes we collected and highlight their potential risks in othering racial minorities. We call for urgent innovations in content moderation policy design and enforcement to address these biases in search outputs.
keywords:
Algorithm audit, algorithmic bias, content moderation, racial bias, search engine
**footnotetext: These authors contributed equally to this work
[…]
Shagun Jhaver, Rutgers University School of Communication & Information, New Brunswick, NJ
1 Introduction
Autocomplete is a technical feature that facilitates the online search process by using the first few entered keywords to predict and offer users a set of query suggestions (Karapapa and Borghi, 2015) (Figure 1). This feature, now ubiquitously available across all major search platforms, social media websites, and online marketplaces, is also available on YouTube, [^1] https://www.youtube.com/ the world’s largest video-sharing platform. Autocomplete improves users’ search experience by reducing typing by about 25 percent on average (Sullivan, 2018). However, the algorithmic ‘nudges’ this feature offers have the potential to induce identity-based biases among YouTube searchers. This is because its autocomplete query suggestions are shaped not only by the inputting user’s previous YouTube activities and location but also by the wider popularity of such queries and the content uploaded by YouTube creators, who often tend to be wealthy and white (Lennard, 2015; Wang et al., 2018). Thus, YouTube autocompletes could incorporate the prevalent biases and stereotypes that some people hold and expose them to other users (Lin et al., 2023).
Such reproduction of stereotypes constitutes representational harm by narrowing demographic groups down to specific traits and exaggerating them (Leidinger and Rogers, 2023). Negative stereotypes about oneself can influence one’s well-being and productivity, whereas negative stereotypes about other groups can permeate public institutions and shape how society and policymakers treat those groups (Levy, 2009). Roy and Ayalon (2020, p.1020) argue that the unsolicited suggestions offered by autocompletion induce “incidental learning,” a “process of unconscious, unplanned absorption of contextual information,” and are likely to influence users’ short- and long-term beliefs and curiosity about different topics. Early investigations of Google autocompletion emphasized that autocompletion stereotypes, viewed and absorbed repeatedly, distort our worldview and perpetuate oppressive social relationships (Cadwalladr, 2016; Noble, 2018).
Figure 1: An example of YouTube’s autocomplete function in action. Typing any keywords (e.g., ”white”) triggers a list of query options that users may select.
Most prior research on autocomplete suggestions centers on algorithmic outputs from general-purpose search engines, especially Google Search (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020). This research focus seems reasonable because Google Search has been the largest global search engine since the early 2000’s. [^2] https://www.britannica.com/topic/Google-Inc However, the use of autocomplete to assist query inputs goes far beyond Google searching. For instance, YouTube, the platform we focus on in this study, has incorporated the autocomplete feature in its search function for many years. Indeed, if we count the number of searches made across all platforms, YouTube emerges as the second-largest search engine after Google (Smith, 2020), offering all its results in a video format. However, YouTube autocomplete has attracted little scholarly attention, perhaps because of YouTube’s public perception as a video-sharing platform rather than a search engine.
Unlike search engines, YouTube allows users to not just view, but also create content on its site (Wotanis and McMillan, 2014). Prior research on YouTube’s participatory culture has characterized it as a critical form of expression, “self-promotion, escapism, and utter play.” (Boxman-Shabtai, 2019) However, it is possible that YouTube’s complex cultural fabric comprising both professional (i.e., traditional media) and amateur (i.e., independent individuals) content production (Boxman-Shabtai, 2019) induces the types of biases that do not appear on search engines.
YouTube algorithms construct and shape content creators’ visibility on the site’s search and front pages as well as in features like autoplaying of ‘related videos’ (Bishop, 2019). However, YouTube releases limited information about these algorithms to maintain a competitive advantage. This black-boxing of algorithmic governance makes accountability difficult (Reynolds and Hallinan, 2024). Thus, there are growing concerns about how cultural producers, who must navigate the increasingly algorithmic nature of content distribution on YouTube, risk losing their visibility and income (Bishop, 2019). Given YouTube’s unique participatory culture and the uneven power relationship between content creators or searchers and the platform, Reynolds and Hallinan (2024) have called for examining YouTube algorithms’ accountability in its specific platform environment. Responding to this call, we focus on YouTube’s search algorithms.
Specifically, we explore what information YouTube users in the United States (U.S.) encounter through the search autocomplete suggestions. Since misinformation or limited information about racial groups, along with media stereotypes, can shape individuals’ perceptions of race (Covington, 2010), we pose the following research question: what information does YouTube provide to users exploring information about racial topics, and how may that information shape users’ understandings of race? We focus on race here to extend prior conversations on the media’s construction of race (Chen, 1996; Squires, 2014).
Importantly, the autocomplete feature not only provides information but may also contribute to the propagation of racial stereotypes. We perform an algorithmic output audit of YouTube autocomplete suggestions regarding four racial groups—Whites, Blacks, Asians, and Hispanics—to surface such stereotypes. We contend that discriminatory race-related outputs generated by algorithms, including autocomplete suggestions, should be interrogated as the most visible signs of deeper issues related to racial biases inherent in sociotechnical systems driving such algorithms.
2 Background and Related Work
2.1 Racial Stereotypes
Since the 1950’s, significant changes in America’s political and social landscape have made racial discrimination illegal and rendered overt racial prejudice socially unacceptable. In response to these broad shifts, social scientists have examined how racial stereotypes have evolved in tandem. Recognizing these changes is crucial in intergroup relations research since shifts in stereotypes are often seen as a necessary prerequisite to reducing prejudice and fostering more positive intergroup interactions.
Since Whites have long constituted the majority of the U.S. population, there exists a cultural tendency toward White ethnocentrism that contributes to shaping Whites positively and other racial groups negatively (Maykovich, 1972). Specifically, US racial stereotypes often manifest as placing Whites and Blacks at the top and bottom of the racial hierarchy, respectively, with other ethnic minorities, such as Asians or Hispanics, in between these extremes (Song, 2004).
Historically, Blacks or African Americans have been negatively stereotyped with respect to a variety of sociocultural characteristics, including their personality, hygiene and criminal conduct. Messages from mass media and political communications often reinforce these stereotypes (Hurwitz and Peffley, 1997). Some scholars maintain that these distorted perceptions of Blacks are rooted in slavery (Lintner, 2004) and Christian mythology (Miller, 1995). Whatever their origin, such racial stereotypes are dangerous because empowered non-minorities can exploit them to justify their privilege and continue to perpetuate harm and discrimination against ‘others.’ For instance, racial segregation in the U.S. was historically based on biased perceptions about Blacks’ personalities, intelligence, and appearance (Lintner, 2004).
Asians or Asian Americans, often described as a ‘model minority,’ are stereotyped as being intelligent, hard-working, and academically and economically accomplished (Lee et al., 2008; Trytten et al., 2012). While this model minority stereotype could be perceived as positive, Lee et al. (2008) argued that it can conceal the diversity among Asians, rouse interracial tensions, and discourage Asians from disclosing their problems and seeking assistance. In addition to the model minority stereotype, Asians have a long history of being marginalized in the U.S. Chen (1996) showed that the physical appearance of Asian male immigrants in the nineteenth century and the work they engaged in were often caricatured to build feminized and infantilized images; the mass media at that time reinforced this image by portraying them as wearing feminine clothes or engaging in housework (Chen, 1996). Further, the U.S. mass media has often portrayed Asian women as exotic, submissive, and quiet (Paner, 2018).
Another US minority, Hispanics, tend to be categorized as a single group despite their heterogeneous and complex ethnic roots (Nelson and Tienda, 2014). They were typically described in early Western films as highly emotional and violent individuals with untidy features, aka the ‘El Bandido’ stereotype (Berg, 2002). This stereotype has gradually evolved into the image of undocumented immigrants, contributing to racist discourses that portray them as criminals who are potentially threatening to American lives and jobs (Pérez Huber and Solorzano, 2015).
Our work extends this prior understanding of racial stereotypes. We examine their current state in the U.S. through the lens of how each racial group is described in YouTube autocomplete algorithm’s suggestions. We aim to identify the perpetuating racial stereotypes reinforced by this culturally influential online platform and highlight recently emerging stereotypes. In doing so, we also bring attention to the algorithmic biases.
2.2 Algorithmic Bias, Search Critiques, and Algorithm Audits
As digital platforms grow, systems designers rely on algorithms to determine what information to present to whom. These algorithms and the interfaces created for them can signal the quality and relevance of different search results and shape people’s perceptions of topics of societal importance (Kay et al., 2015). However, both algorithms and interfaces can exhibit biases in their representation of information. While there is no universally agreed-upon definition of algorithmic bias, researchers from various disciplines have addressed its different manifestations. In social science, scholars tend to focus on how algorithmic bias perpetuates social discrimination and affects social equity.
The research presented here can be placed within the broader critique of algorithmic biases in content moderation, though we focus on search feature outputs rather than AI-assisted regulation of potentially inappropriate social media posts (Jhaver et al., 2019). Specifically, this article aligns with approaches that aim to assess the extent to which search outputs are: problematic; characterized by offensive, stereotypical, inappropriate, and discriminatory terms; and/or racist (Rogers, 2023). Our method of probing YouTube’s search feature with specific queries could be positioned alongside approaches like algorithm audits (Sandvig et al., 2014) and ethical hacking practices that search for vulnerabilities (Roberts, 2019).
Prior research has shown algorithmic biases in the operation of search engines regarding what they index, what they present to specific users, and—most relevant to our work—how their autocomplete suggestions reinforce social discrimination (Kay et al., 2015). For example, Baker and Potts (2013) showed that Google search produces a higher number of negatively stereotyped autocomplete suggestions for searchers with Muslim, Jewish, Gay, or Black identities. Lin et al. (2023) collected autocomplete predictions from three leading search engines—Google, Bing, and Baidu—and showed disparities in their toxicity scores with respect to gender, race, and sexual orientation.  Leidinger and Rogers (2023) showed that autocompletes from Yahoo, another general-purpose search engine, include a large number of negative stereotypes for Latinos, e.g., portraying them as stupid and loud.
Our work contributes to this line of research by deploying a novel methodological approach and conceptual focus to understanding the prevalence of racial biases in YouTube search autocompletes. While prior studies have primarily focused on only the text of autocomplete suggestions (Al-Abbas et al., 2020; Baker and Potts, 2013; Leidinger and Rogers, 2023; Roy and Ayalon, 2020), we analyze autocomplete texts in conjunction with the video results for each suggestion. This helps us present a more nuanced understanding of how YouTube users’ interactions with its search feature may influence their perspectives. Our work is the first to deploy critical discourse analysis (CDA) in addition to inductive analysis to uncover not only how language and discourse practices of autocompletes reflect problematic socio-cultural norms but also how they reinforce hegemonic whiteness and racist ideologies.
It is possible that biased queries entered by users, in turn, could be influencing the autocomplete algorithm to produce stereotypical outputs. However, given the black-box nature of autocomplete algorithm, the extent to which this happens remains unclear. Some recent evidence suggests that the impact of search volume on autocomplete suggestions may be lower than is generally assumed and, in some instances, prior user queries may not feature at all in certain suggestions (Graham, 2022). Thus, we largely focus in this article on the audit of autocomplete outputs rather than on inferring how biased user queries could influence the production of stereotypes in autocompletes.
3 Methods
3.1 Development of Input Search Queries
We began our inquiry by selecting the following four categories of racial groups: “White,” “Black,” “Asian,” and “Hispanic.” This selection was based on racial categories most frequently included in prior research on stereotypes or biases in search engine autocompletes and language models (Kirk et al., 2021; Leidinger and Rogers, 2023). Next, we developed a series of input queries based on combinations of group indicators and verb phrases. We leveraged the synthetic patterns of input queries employed in previous autocomplete studies (Al-Abbas et al., 2020; Baker and Potts, 2013; Roy and Ayalon, 2020) to guide this query creation. We combined the following verbs—is, are, does, do, can, should—with racial group indicators in the form of declarative and interrogative sentences. Table 1 shows the group terms and synthetic patterns we used to structure our input queries. In total, we curated 84 input queries for each of the four racial groups through this process.
Category Group terms Synthetic patterns
White White/Whites, White person/people, White man/men, White boy/boys, White woman/women, White girl/girls Is a(an)/are [Group Terms]
Does a(an)/Do [Group Terms]
Can a(an)/Can [Group Terms]
Should a(an)/Should [Group Terms]
A(An) [Group Terms] is/are
A(An) [Group Terms] can/can
A(An) [Group Terms] should/should
Black Black/Blacks, Black person/people, Black man/men, Black boy/boys, Black woman/women, Black girl/girls
Asian Asian/Asians, Asian person/people, Asian man/men, Asian boy/boys, Asian woman/women, Asian girl/girls
Hispanic Hispanic/Hispanics, Hispanic person/people, Hispanic man/men, Hispanic boy/boys, Hispanic woman/women, Hispanic girl/girls
Table 1: The structure of input queries used to guide our data collection.
3.2 Data Collection
We collected data from July through August 2023 by gathering YouTube autocomplete suggestions for our input queries. Figure 2 shows an example of a YouTube search with a sample input query and its autocompletes; we collected such autocompletes for each query. We automated our data collection by developing Python scripts that simulated YouTube search querying actions. This automated process comprised three stages: 1) accessing the YouTube.com website, 2) entering an input query in the YouTube search box, and 3) collecting autocomplete results, if available, and storing them. We used Chrome as a web browser and accessed YouTube in incognito mode without logging in to any account to prevent the effects of personalization and browsing history. Our scripts opened and closed a new browser window for each query to ensure that the search history of other input queries did not influence the autocomplete suggestions for any single query.
Figure 2: Screenshot of a YouTube search for one of our input queries. It shows five of the autocomplete results we collected that correspond to the query.
We used a virtual private network to collect data from multiple locations in the U.S. to account for location-specific biases (Ribeiro et al., 2020). In particular, we collected YouTube autocompletes from five states: New Jersey, California, Georgia, Texas, and Washington. We conducted our data collection three times, each separated by a few days, to account for the temporal variations in algorithmic suggestions. Through these efforts, we attempted to build a geographically representative and temporally stable dataset of autocomplete suggestions across the U.S.
Analyzing the data samples from three different waves and five different states, we found minimal differences among them. We integrated these samples and removed duplicate results to create a combined dataset for further inquiry. While this step precluded prioritizing more geographically and temporally stable autocompletes, it allowed us to work with a more comprehensive set of suggestions. Examining the resulting dataset, we also noticed that some autocomplete suggestions were not related to the racial groups. Thus, we further removed the irrelevant autocomplete suggestions based on whether the autocomplete refers to non-human rather than non-human subjects. After removing duplicates, our dataset consisted of a total of 241 unique autocompletes.
We describe more details of our data preparation process and descriptive statistics in the Supplemental Material.
3.3 Data Analysis
We used inductive analysis (Elo and Kyngäs, 2008) to examine our data and extract concepts and categories that describe the underlying phenomena. Two authors conducted a bottom-up open coding process on the platform ‘Dedoose’ which allowed the authors to collaborate in real-time.
We began by developing a situated understanding of each autocomplete suggestion by examining about ten videos displayed as top search results when selecting that suggestion. This examination helped us recognize the contextual and cultural meaning of autocomplete suggestions and we leveraged this nuanced understanding to develop our open codes. This careful, context-driven coding process was important because some of the autocomplete results were closely related to memes or video content in their corresponding search results. For instance, the autocomplete “when an asian is kidnapped” was related to comedy content created by an Asian YouTuber. Thus, it was important to search for and understand the situated meanings of autocompletes rather than interpret them in isolation.
Then, we developed the common themes arising from our open codes and their corresponding autocomplete data by grouping similar codes into conceptual categories. At this point in the process, when any pre-existing categories seemed not to directly align with an emergent theme, we carefully considered whether to introduce a new category, revise an existing category, and/or create a subcategory. This process was iterative and we compared our emerging categories with one another and with the underlying data at each stage of our analysis. Some autocompletes represented miscellaneous sociocultural contexts that did not warrant inclusion in or creation of a separate category; we excluded such autocompletes from further analysis. In total, we grouped 217 autocompletes into a two-tier category structure.
Throughout the process of qualitative coding and categorization of autocomplete results, we incorporated critical discourse analysis (CDA). CDA is a systematic approach to examining a text, event, or discursive practice to identify how it originates from and is affected by power relations (Fairclough, 1993). We employed this framework to analyze each autocomplete as concurrently a text (language produced in a discursive event), a discourse practice (the production, distribution, and consumption of a text), and a sociocultural practice (hegemonic struggles that ideologically shape a discursive practice) (Fairclough, 1995). This meant that during data analysis, our focus was less on examining the lexical meanings of each autocomplete suggestion than on the hegemonic function the suggestion performed or indicated within the sociocultural practices represented by our five main categories. Using this approach, we aimed to identify various facets of each discursive event, thereby revealing particular power dynamics reflected in that event. We engaged in regular discussions about how the autocomplete suggestions in our data reproduce and legitimize social inequalities, and our coding sought to identify various forms of domination and biases.
4 Results
Our analysis generated five main categories representing distinct contexts in which we observed racial biases: Appearance, Ability, Culture, Social Equity, and Manner. Under these five main categories, we identified a total of 13 subcategories (Table 3).
Main category Subcategory Description % (n)
Appearance Personal Hygiene The activity of cleaning the body and maintaining a good appearance 3 (6)
Skin Tone The skin color originating from racially different genetic factors 5 (10)
Ability Talent One’s abilities for artistic or physical performances 22 (47)
Financial One’s abilities to purchase goods or services 2 (4)
Intellectual One’s abilities to learn and study 2 (4)
Culture Cultural Heritage Conventional cultural elements of one society with a shared identity 21 (45)
Ethnic Humor The humor based on racial, ethnic, or national stereotypes 13 (28)
Language How someone communicates verbally 6 (12)
Relationships How people interact and connect with each other 9 (19)
Social Equity Diversity/Inclusion Practices of embracing socially diverse groups of people 2 (5)
Racial Justice Pursuit of fighting against injustice toward a certain racial identity 10 (22)
Manner Aggression Antisocial and violent traits 5 (11)
Inappropriate Behavior Being idle or behaving in a socially inept way 2 (4)
Table 2: Categories of racial stereotypes that emerged in YouTube autocomplete suggestions. % and n refer to the percentage and number of autocompletes belonging to each subcategory in the analyzed data, respectively.
4.1 Appearance
Personal Hygiene. This category is notable because only Blacks were subject to autocompletes related to personal hygiene. These autocompletes referenced Black people washing their hair or beard (“how often should a black man wash his hair,” “how often should a black woman wash her hair,” “how often should a black man wash his beard”) and shaving (“should a black man shave everyday”). Figure 3 shows the preview of a video resulting from the suggested autocompletes, which contains advice for preventing white flakes in hair. [^3] https://www.youtube.com/watch?v=CZCuI4x1sUc Another autocomplete asked, “do black people get lice.” It is possible that these autocompletes reflect highly sought-after searches by Black people for the information needed to take care of textured hair. However, these questions are raised only for the Black group, raising the possibility that they reflect the racial stereotypes of Black people having poor hygiene. Historically, a deep-rooted racial stereotype maintained that dark or non-white skin was associated with uncleanness (Brown, 2009); indeed, this stereotype served as a rationale for racial segregation in the U.S. (Lintner, 2004). These YouTube autocompletes suggest that the negative stereotypes about Blacks’ hygiene unfortunately persist.
Figure 3: A YouTube video suggested by the autocomplete “how often should a black man wash his hair.”
Skin Tone. Similar to ‘Personal Hygiene,’ Blacks were the main subjects of autocompletes in the ‘Skin Tone’ category. Autocompletes related to skin tones included “can a black person become fair.” These bleaching-related autocompletes reflect both an interest in lighter skin and the implicit negative view of darker skin (Hunter, 2011). We observed no autocompletes related to the action of changing natural skin tones for other races. Other autocompletes show how Black skin is perceived as unique and distinct from other races. Examples include autocompletes asking about the need for Black people’s skin protection (“should black people wear sunscreen,” “do black people get sunburn”) and changes in their natural skin tones (“can black people blush,” “can black people tan”). These results suggest both an otherization and a general lack of awareness about Black skin characteristics.
4.2 Ability
Talent. Our analysis revealed racial stereotypes associated with talent in music, dance, and sports that mainly addressed White and Black groups. For White groups, the autocompletes related to their artistic talent included both positive (“white girl can rap,” “white girl can sing”) and negative (“white people can’t dance”) elements. Regarding physical talent, several autocompletes reflect stereotyped perceptions that Whites are not as athletic as Blacks (“white men can’t jump,” “white boys can’t dance”). Indeed, the videos that such autocompletes guided users to related to a 1992 sports comedy movie called “White Men Can’t Jump,” which leverages the widely held stereotype that Black race is a positive indicator of physical ability (Felson, 1981). Autocomplete suggestions linked to this movie’s title decades after its release date suggest a cultural appetite to mock or stereotype White athletes as being inferior in their physical performance.
Similarly, most autocompletes for Blacks referenced their artistic or physical ability in both positive (“black people can dance to anything,” “black man can sing”) and negative (“black man can’t dance,” “black man can’t play basketball”) terms. Notably, we found some autocompletes related to Blacks’ exceptional athletic abilities (“are black people more athletic”), which aligns with the previously reported stereotype (Biernat and Manis, 1994) that Blacks are athletically superior to Whites. Those who did not meet such expectations were negatively stereotyped. For example, the autocompletes “black man can’t dance,” or “black man can’t play basketball” reflect a stereotype that Black people should be good at dancing or basketball naturally, and the resulting videos mock Blacks who are not good dancers or sportsmen. In sum, these findings show that YouTube autocompletes reflect racial stereotypes based on widely held beliefs about the relationships between race and physical abilities.
Financial. Only Blacks were subjects of autocompletes that belonged to financial ability. All autocompletes here were associated with a lack of financial resources and acumen, which reinforces poor Blacks as the representative public image of poverty. Interestingly, this is consistent with how network TV news and weekly news magazines have historically grossly overrepresented Blacks in their portrayal of poor people as a whole (Gilens, 1996). Autocompletes in this category promoted the stereotype that Blacks cannot purchase or own expensive products, such as fancy cars or clothes (“black man cannot own g wagon,” “black woman can’t buy a dress,” “a black man can’t have a suitcase”).
While these results superficially reinforce the long-standing stereotypes of Blacks as poor, some video results for these autocompletes counter this stereotype. For example, the top video result for “black woman can’t buy a dress” portrays a White shop assistant and manager treating a Black woman attempting to buy a bridal gown at a shop differently than they treat a White woman buying a gown (Figure 4). [^4] https://www.youtube.com/watch?v=XMcRS2-la3E This video undermines the stereotypical attitudes toward Black people regarding their financial ability (it also critiques the assumptions of criminality and poor hygiene associated with Blacks). Similarly, resulting videos for other autocompletes about financial ability dismiss the assumption that Black people are poor by explicitly showing that Black people can and do own expensive products. This shows that in some cases, users who review and watch the search results may get exposed to video content that helps diminish negative stereotypes of Blacks regarding financial ability.
Figure 4: A YouTube video suggested by the autocomplete “black woman can’t buy a dress.”
Intellectual. Intelligence-related autocompletes appeared for Asians and Blacks. In the U.S., Asians have historically been perceived as intelligent, hard-working, and pursuing high educational achievements (Trytten et al., 2012). Autocompletes such as “are asians smarter” reflect this model minority stereotype. Unlike Asians, Blacks have often been negatively stereotyped regarding their intellectual and academic abilities (Lintner, 2004; Miller, 1995). In line with this, we found autocompletes regarding Blacks’ intelligence: “black people can’t name one african country,” and “black man can’t say beginning.” Videos for the first autocomplete show the lack of knowledge of some Black people about Africa, which suggests their ignorance of Blacks’ historical or ethnic roots. The second leads to a video of a South African President, who is Black, repeatedly mispronouncing his words when attempting to say “in the beginning.” These cases show how videos uploaded to YouTube position individual mishaps to denigrate racial groups under the guise of humor. YouTube’s automated use of such content to build autocomplete suggestions further removes necessary context and can thus perpetuate racial stereotypes.
4.3 Culture
Cultural Heritage. Our analysis reveals that YouTube autocompletes reflect long-standing debates about cultural appropriation, described as “the use of one culture’s symbols, artifacts, genres, rituals, or technologies by members of another culture” (Rogers, 2006, p.476). We found that concerns about such appropriation appear in YouTube autocompletes related to several aspects of cultural heritage, such as cultural practices and culturally specific attire.
First, we found autocompletes that indicated appropriation of cultural practices, highlighting asymmetric power relations between racial groups. For example, several autocompletes related to White groups coopted indigenous Black cultures while belittling its native contexts (e.g., “white girl can dance african”). Other autocompletes revealed how White people could appropriate specific aspects of popular Asian culture, e.g., “can a white guy won [sic] an asian beauty pageant,” “can a white person become a kpop idol.” The keyword ‘Black’ elicited one autocomplete related to the adoption of Asian cultural elements (“can a black person become a kpop idol”).
Second, autocompletes associated with culturally specific attire and beauty standards also indicated instances of cultural appropriation, where members of the majority group coopted traditional garments, hairstyles, or makeup practices of other minority groups. For example, autocompletes related to garments or hairstyles for White groups included traditional styles with deep origins in Asian or Black culture, e.g., “can a white person wear a kimono,” “can a white person wear a durag,” “can white girls get braids,” and “can a white boy get waves” (Figure 5). [^5] https://www.youtube.com/watch?v=gSZ8U64EL-Q Additionally, we found instances of blackfishing, i.e., the practices of modifying one’s appearances physically or digitally or using terms and language patterns linked to Ebonics for racial commodification (Stevens, 2021, p.1). For example, we found autocompletes for White groups that included references to the makeup style of White people attempting to look like Blacks (“the white woman who turned black”).
Figure 5: A YouTube video suggested by the autocomplete “can a white person wear a durag.”
In contrast, for Asians and Blacks, autocompletes related to cultural attire rarely reflected acts of cultural appropriation toward other cultures. Instead, the autocompletes for Blacks were related to changing their natural hairstyle (“how can a black person get curly hair,” “can a black man straighten his hair”) or achieving a popular hairstyle or makeup (“how to do a black person’s hair,” “how to do a black person’s makeup,” “how to do a black man bun”).
The autocompletes for Asians often included the term “Asian Baby Girl (ABG)” (“what is an asian baby girl,” “what is an abg asian baby”). Asian Baby Girl is a slang term referring to “a young Asian woman who displays stereotypical traits, such as enjoying clubbing, wearing excessive makeup and tattoos, drinking bubble tea, wearing revealing clothes, etc” (Wiktionary, 2023). Though the popularity of such language can subvert stereotypical images of Asian women as submissive or quiet, it also associates them with negative attributes, like aggression and violence (Wu, 2023). Thus, autocompletes about ABG could increase awareness about the diversity of Asian women but, in doing so, risk generating new negative stereotypes about them.
In sum, our analysis reveals that YouTube autocomplete suggestions tend to be unidirectional, i.e., supporting ways that dominant racial groups (usually White) appropriate the cultural heritage of racial minorities but not vice versa.
Figure 6: A YouTube video suggested by the autocomplete “when an asian is bad at maths.”
Ethnic Humor. Schutz defined ethnic humor as “humor directed at racial and nationality groups, denigrating alleged attributes of those groups” (Schutz, 1989, p.167). Ethnic stereotypes or shared beliefs toward specific ethnic groups are the basis of such humor (Boxman-Shabtai and Shifman, 2015). In our analysis, Asians have the largest number of autocompletes in this category. Most of these autocompletes led to content created by the Chinese-Irish comedian Steven He O’Byrne (12.3 million subscribers). [^6] https://www.youtube.com/@StevenHe O’Byrne’s videos usually parody Asian dads and mock East-Asian parenting stereotypes. For instance, the video content corresponding to the autocomplete “when an asian is bad at maths” shows O’Byrne portraying Asian parents who have a high expectation for their children’s academic achievements and are harsh toward them when grades do not meet their standards (Figure 6). [^7] https://www.youtube.com/watch?v=6lMpjmPiTWE Other autocompletes, like “when an asian is kidnapped,” “when an asian is president,” and “when an asian is your substitute,” lead to videos with similar stereotypical images of Asians. While these videos presumably aim to generate clicks and their associated revenue through humor, they reproduce stereotypes by portraying Asians as alien or different and endorsing the ‘model minority’ myth, which characterizes Asian Americans as being financially and educationally more advanced than other minority groups (Lee et al., 2008).
Language. Language attributes (such as accents) tend to be understood as a strong indicator of ethnic categorization (Rakić et al., 2011). We found several autocompletes related to accents or speaking styles of each racial group. For example, the autocompletes for speaking styles included, “how to do a white girl voice,” “how to do a white person voice,” and “how to do a black man voice.” Autocompletes for accents referenced the unique mode of pronunciation of certain racial groups (“black man can’t pronounce,” “what does an asian accent sound like,” “all latino accents,” “woman with spanish accent”). Videos resulting from such autocompletes often depicted individuals mimicking or mocking the voices and accents of other racial groups. This suggests that the multicultural aspects of U.S. languages are still being leveraged to otherize specific groups.
Relationships. Many autocompletes in this category highlight anxieties about interracial relationships. For example, some autocompletes reflect curiosity or potential worries about relationships between White women and people with darker skin (“do white girls like brown guys,” “do white girls like latinos”). Similarly, many autocompletes referenced relationships between Black people and women from different cultures (“can a black man marry an arab woman”). These results often portray interracial relations as incongruous and reflect an unfavorable view of darker skin in the relationship contexts.
The most severe tension was found regarding the interracial relationships or marriages between White and Black individuals (“can’t date a black woman until my grandpa died,” “black people can’t marry meme,” “black people can’t marry white people”). These results indicate that cultural remnants of the historical prohibitions of interracial marriage and fears around miscegenation persist (Pascoe, 1991). There was only one result for Asians in this category: “asian men are bitter.” Though seemingly unrelated to romantic relationships, the top resulting videos of this autocomplete had titles such as “Asian men are least desired.” These videos depicted the struggles of Asian men with dating and pursuing romantic relationships, thereby perpetuating their historic feminization and infantilization (Chen, 1996).
4.4 Social Equity
Diversity/Inclusion. Whites, Blacks, and Hispanics had autocompletes associated with the issues of diversity and inclusion. Curiously, most of these autocompletes referred to the underrepresentation of racial minorities in popular culture (“can a black man be captain america,” “is there a hispanic disney princess” Figure 7). [^8] https://www.youtube.com/watch?v=AQLUeshn3nE This reflects a cultural appetite for seeing racial diversity and encouraging inclusive casting decisions in mainstream media. While whitewashing has historically been salient in the entertainment industry, the percentage of White characters depicted on-screen has slightly declined in recent years, while the depiction of Asian characters has significantly increased (Smith et al., 2023). However, no notable changes have occurred in the prevalence of Black or Hispanic characters over the same period (Smith et al., 2023). Seemingly in response to these trends, YouTube autocompletes in this category reveal frustrations about persistent inequalities in the racial representation of Black and Hispanic character portrayals.
Figure 7: A YouTube video suggested by the autocomplete “is there a hispanic disney princess.”
Racial Justice. We observed autocompletes in this category only for the White and Black racial groups. First, we found autocompletes that referenced ‘White Lives Matter,’ a racist counter-movement that emerged in response to the ‘Black Lives Matter’ movement (“do white lives matter one minute,” “do white lives matter shorts”) (Figure 8). [^9] https://www.youtube.com/watch?v=GlMYUFDb8qg ‘White Lives Matter’ uses a variation of the Black Lives Matter core slogan to weaken the latter’s message about addressing unwarranted police violence against Blacks (Goodman et al., 2023). The slogan ‘White Lives Matter’ appeared verbatim in YouTube autocompletes. Despite video results for the query ‘White Lives Matter’ displaying diverse voices and criticisms on contested topics, text suggestions of YouTube autocomplete itself were solely generated based on the messages of ‘White Lives Matter,’ and no analogous autocompletes were associated with ‘Black Lives Matter’ messaging. This indicates that the autocompletes curated by YouTube do not always balance perspectives from opposing sides of an issue and, in this case, give short shrift to the side combatting racial disparities.
One autocomplete indirectly reflected the perspective of Blacks on racial inequality: “can’t keep a black man down,” a lyric excerpt from 2Pac’s song “Trapped,” which addresses police brutality and harassment faced by African Americans (Gaines, 2022). This autocomplete and its resulting videos highlighted ongoing racial inequality and police brutality toward African Americans in the U.S.
Figure 8: A YouTube video suggested by the autocomplete “do white lives matter one minute.”
4.5 Manner
Aggression. Our results reference perceptions of Blacks and Asians as being aggressive. Blacks had the highest number of autocompletes, which often described them as dangerous (“pretend there is a black man chasing you,” “run like a black man is chasing you”) or criminal (“a black woman is accused of shoplifting”). This aligns with long-held negative stereotypes of Black violence and aggression (Gilens, 1996). We also found one autocomplete that suggests a desire to circumvent this stereotype: “every black man should have a latte.” While this does not relate to aggressive traits at face value, the top resulting video shows a Black stand-up comedian poking fun at the violence stereotype by recommending Black men hold a latte to appear ‘safe’ to law enforcement officials. Additionally, we found aggression-related autocompletes for Asians: “asian man goes crazy.” The top resulting video for this shows an Asian male game streamer screaming violently while playing games. Other videos show Asian men engaging in gun violence. While such aggressive images are typically not associated with Asian men in the U.S., these videos depict Asian men as not always docile and potentially violent.
Inappropriate Behavior. This category contains autocompletes that describe Black people as lazy or behaving inappropriately in social settings. One longstanding negative stereotype of Black people maintains that they are lazy, which is also posed as an explanation for other undesirable attributes, such as their disproportionate poverty (Gilens, 1996). The autocomplete “black people can’t hear smoke detectors” depicts this stereotype; it led to videos that show Black people as being too lazy to change the low battery of smoke detectors even though their alarms kept sounding. Figure 9 presents a series of YouTube shorts that show how creators promote this stereotype.
Several autocompletes describe Black people as lacking the self-control needed to behave appropriately in social settings. For instance, this category included “black man cant [sic] hold his laugh,” and “black man can’t stop laughing.” The resulting videos show instances where a Black man cannot refrain from laughing in a serious situation where laughter is clearly inappropriate.
Figure 9: A series of YouTube shorts suggested by the autocomplete “black people can’t hear smoke detectors.”
5 Discussion
We began this study with the goal of examining the information that YouTube’s search feature provides to users seeking video content on race-related topics. We found that besides listing offerings that assuage users’ curiosities on these topics, YouTube plays a role in spreading racial stereotypes through its autocomplete function, a feature that incidentally cannot be turned off.
Our analysis reveals that race-related autocompletes embody troubling stereotypes regarding Blacks’ personal hygiene, financial abilities, language use, aggression, and laziness; Whites’ athletic abilities; and Asians’ image as the high-achieving model minority. Notably, we observed that YouTube search amplifies distorted racial views that appear in mainstream entertainment media by reproducing titles or dialogues from existing media content (e.g., “white men can’t jump”). We posit that this can be attributed to YouTube’s participatory culture (Khan and Malik, 2022), where users actively contribute to content creation aimed at capturing popular attention and resulting in monetary gains. Thus, an interplay of mainstream media and amateur online cultures could be perpetuating racial stereotypes on the platform. Creators’ communal and commercial aspirations (Boxman-Shabtai, 2019) could also account for our finding that racial stereotypes on YouTube are often subtly delivered in the form of ethnic humor and memes. These genres and formats may make naive viewers embrace group stereotyping as a new and enjoyable form of entertainment rather than recognize its potential harm toward racial minorities, as evidenced by the millions of views and likes such content generates.
In addition to reinforcing racial stereotypes, our analysis of YouTube autocompletes reflects tensions about interracial relationships, and Blacks’ and Hispanics’ concerns about their media representation. The norms of whiteness appear to prevail, with racial minorities struggling to combat structural inequalities (“can a black man be captain america,” “is there a hispanic disney princess”). These results serve as a lens through which to examine the current state of race in the U.S. Though some popular narratives maintain that we live in a ‘post-race’ society (Squires, 2014), the evidence of racial tensions we found in our analysis indicates otherwise. Despite a range of efforts to reduce racial discrimination, our study indicates the need for additional initiatives to promote racial harmony and resist the othering of racial minorities.
Our results highlight ethical concerns about the potential ‘nudge effects’ of autocomplete suggestions on YouTube users. Graham (2023) noted that autocompletes can be biased in ways that might not be clear on an individual level but become apparent when evaluating a larger sample. Our analysis reveals evidence of similar aggregated discrimination in race-related queries on YouTube. For example, we find that while Black struggles to combat oppression are diminished, e.g., through the removal of references to “Black Lives Matter” in autocomplete suggestions, “White Lives Matter” continues to feature in these suggestions despite its known associations with hate groups. We also found traces of power imbalance through various ways of cultural appropriation reflected in YouTube autocompletes, showing how asymmetric racial values are embedded in YouTube’s technological configurations. Our unique methodological approach has helped us surface these biases, and it can also inform future analyses of autocomplete outputs, especially on social media and video-sharing platforms.
Prior studies have shown how biased autocomplete suggestions can impact users’ votes in political elections (Epstein et al., 2024), lead to hiring discrimination (Lin et al., 2024), and shift attitudes toward societal issues, e.g., support for fracking and legalizing the death penalty (Williams-Ceci et al., 2024). Thus, there is cause for concern that repeated exposure to racial stereotypes in YouTube autocompletes may adversely influence users’ attitudes and behaviors toward different racial groups. The black-box nature of such AI-based technologies is likely to prevent users from gaining access to information about how outputs were generated and how they could embed identity-based biases (Rai, 2020). In fact, users may be totally unaware of AI biases and fail to notice the influence of these biases on their attitudes (Williams-Ceci et al., 2024). Future work can build upon this research by analyzing the cognitive processes that mediate how YouTube users are influenced by biases in autogenerated content on the site.
Indeed, previous research on Google Search shows that its autocomplete outputs are sourced not just from web content and users’ search activity but also from external digitization projects, especially Google Books (Graham, 2023). Lack of information about how autocompletes are produced allows platforms to avoid taking the full ethical responsibility for problematic suggestions, even in cases when they result from machine learning biases rather than prior search queries. Therefore, we call for greater transparency regarding autocomplete sources and algorithms to facilitate a more precise attribution of problematic suggestions.
Our findings indicate the need for moderation strategies for the ethical deployment of AI-based tools such as autocompletes. Prior research on moderation in search engine autocompletes has shown that platforms moderate derogatory stereotypes about various demographic categories (Leidinger and Rogers, 2023). Google even has an autocompletion moderation policy to prevent autocompletes “that might be violent, sexually-explicit, hateful, disparaging or dangerous” (Sullivan, 2020). Yet, as our findings show, YouTube, owned by the same parent company as Google, perpetuates racial stereotypes, highlighting the need for improved policy and corporate-wide enforcement. As Rogers (2023) points out, platforms like Google should take actions that exceed mere ‘quiet patches’ of infamous autocompletions highlighted by journalists and researchers on English language sites by ensuring that adequate remedies occur for all their products and language offerings.
This study has some limitations. We observed an under-representation of Hispanics in our analysis of autocompletes. However, it is possible that this result is influenced by other terms referring to the “Hispanic” race (e.g., “Latino,” “Mexican”) being used colloquially more often than similar counterparts (e.g., “Caucasian,” “African-American”) for other races. Future research on autocomplete audits should also collect queries using such colloquial terms.
A Black person can
can’t date a black woman until my grandpa died Does a Black woman
A Black woman can
can’t keep a black man down Does a Black man
Can a Black man
how often should a black man wash his hair Should a Black
Should a Black man
A Black man should
how often should a black woman wash her hair Should a Black
Should a Black woman
A Black woman should
how often should a black man wash his beard Should a Black man
black woman can’t buy a dress A Black woman can
a black man can’t have a suitcase A Black man can
run like a black man is chasing you A Black man is
every black man should have a latte A Black man should
black man can’t pronounce Black men can
black man can’t dance Black men can
black man can’t play basketball Black men can
black man can’t say beginning Black men can
black man cannot own g wagon Black men can
black man can’t stop laughing Black men can
black man cant [sic] hold his laugh Black men can
black people can’t marry meme Black people can
black people can’t marry white people Black people can
black people can’t name one african country Black people can
black people can’t hear smoke detectors Black people can
Asian what does an asian accent sound like Does an Asian
what is an asian baby girl Is an Asian
what is an abg asian baby Is an Asian girl
when an asian is bad at maths An Asian is
when an asian is kidnapped An Asian is
when an asian is president An Asian is
when an asian is your substitute An Asian is
asian man goes crazy An Asian man is
Hispanic is there a hispanic disney princess Is a Hispanic
all latino accents A Hispanic should
woman with spanish accent A Hispanic girl should
Table 4: This table lists autocomplete suggestions and the input queries they derived from for cases where the suggestion did not contain the input query. This highlights that YouTube’s autocomplete generation does not strictly follow the exact syntactic patterns of input queries.